OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms?

OpenStudy (anonymous):

options?

OpenStudy (anonymous):

A. -1 + 6x = 6 B. 10x - 5 = 6 C. 4x + 1 = 6 D. 4x - 5 = 0

OpenStudy (anonymous):

4x-5+6x=6 10x-5=6 10x=11

OpenStudy (anonymous):

B

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@Maccrea dont just give an answer please we are suppose to help them learn not do homework for them

OpenStudy (anonymous):

You should try to work this out before you post it on here

OpenStudy (anonymous):

i Did..

OpenStudy (anonymous):

Well I don't have a chance to since you do that for me, I know how to use the site